# Towards Multi-Granularity Memory Association and Selection for Long-Term Conversational Agents

### v1

@@ 1 Introduction
Large Language Models (LLMs) have showcased remarkable conversational abilities, enabling them to serve as personalized assistants [22, 47, 23] and handle a wide range of applications, such as customer service [34] and software development [36]. However, as user-agent interactions increase, the volume of conversational history grows significantly. Despite their impressive conversational abilities, LLMs struggle with maintaining long-term conversational memory due to their limited context length [27], which makes it challenging to retain a comprehensive record of user interactions and preferences over time. This limitation undermines their ability to generate coherent and personalized responses [50, 33]. The development of retrieval-based external (non-parametric) memory systems [32, 35] has emerged as a promising solution. By storing interaction histories and retrieving relevant information when needed, memory systems enable LLMs to recall user-specific details from past dialogues and deliver more tailored responses.
@@
Recent studies on external memory systems of LLM agents primarily depend on the Retrieval-Augmented Generation (RAG) pipeline and have explored diverse aspects of memory segmentation and construction for effective retrieval [51, 56]. For memory segmentation, existing approaches mainly adopt single-granularity to segment conversations. They utilize session-level chunks as retrieval units [28, 26, 45], while others employ finer-grained turn-level segmentation to capture details [50, 54]. Recent advances introduce topic-aware segmentation techniques [33, 44] that group dialogue based on semantic coherence, enhancing topic-consistent retrieval. Additionally, several works [49, 45, 59] generate memory summary, condensing key information into compact representations to improve retrieval efficiency. Regarding memory construction, researchers have explored structured organization paradigms to enhance long-term knowledge retention [25, 6, 7]. Methods like RAPTOR [42] and MemTree [39] employ hierarchical tree structures to encode multi-scale relations between memory units. Inspired by cognitive mechanisms [46], HippoRAG [8, 9] implement graph-based memory architectures that simulate neural memory consolidation processes, establishing relations between entities.
@@
Figure 1: An example of the multi-session query. The agent needs to synthesize information from both Conversation 1 and Conversation 2 to answer the question. The red underlines highlight information directly relevant to the answer, while the bold text emphasizes the same details between conversations. By generating multi-granularities (e.g., summary and keyword), the agent can better establish connections between conversations, enhancing full memory retrieval. Additionally, chart (a) illustrates the performance across different granularities, where ‘optimal granularity’ means the result using best-suited granularity for each query.
@@
However, despite their progress, current approaches exhibit two critical limits: i) Insufficient Multi-Granularity Memory Connection. While existing methods endeavor to organize memories into topological structures (e.g., knowledge graphs [8, 9] or trees [42, 39]), they predominantly concentrate on singular granularity levels—either entities or session summaries. This single-scale paradigm fails to model cross-granular interactions between memory units, resulting in retrieving only partial useful information. As demonstrated in Figure 1, answering multi-session queries requires establishing semantic links across granularities (e.g., connecting Conversation 1 and Conversation 2 through shared keyword/summary). Failure to establish links results in partial retrieval (e.g., retrieving only Conversation 1), which leads to incorrect answers. ii) Lack of Adaptive Multi-Granular Memory Selection. Current methods mainly rely on fixed granularity strategies (e.g., session/turn segmentation or LLM-generated summaries [49, 59, 50]), which often lead to incomplete context recall or noise due to improper granularity. Although topic-aware segmentation [33, 44] enhances intra-chunk coherence, they lack adaptive mechanisms to select optimal granularity per query. Our empirical analysis in Figure 1(a) reveals that adaptively choosing optimal granularity per query (e.g., balancing noise reduction in summaries/keywords with information retention in raw sessions) yields substantial performance gains. This highlights the necessity of granularity selection to resolve the inherent noise-information trade-off.
@@ 2 Methodology
This section defines the task and data format, constructs a dynamical memory association framework, details an entropy-based router for optimal granularity, and outlines strategies for retrieving and filtering high-quality contextual information for response generation.
@@ 2.2 Multi-Granularity Association Construction
Existing methods mainly encode memories into vector libraries (e.g., session-level chunks) and directly retrieve information via similarity search [13, 19, 28]. However, such approaches overlook deeper associations between memories. To address this, we propose an associative memory construction process that captures multi-granular relationships.
@@
Dynamical Memory Association. When a new memory $\mathcal{M}_{\text{new}}$ is added, the current memory bank $\mathcal{M}_{\text{cur}}$ is updated as $\mathcal{M}_{\text{cur}}\leftarrow\mathcal{M}_{\text{cur}}\cup\mathcal{M}_{\text{new}}$. To establish associations between $\mathcal{M}_{\text{new}}$ and historical memories, a Gaussian Mixture Model (GMM)-based clustering strategy [38] is used. All memories are encoded into dense vectors $e(\mathcal{M}_{\text{cur}})$ and $e(\mathcal{M}_{\text{new}})$ by Contriever [13], and pairwise similarity scores are computed between $e(\mathcal{M}_{\text{new}})$ and each $e(\mathcal{M}_{j})\in e(\mathcal{M}_{\text{cur}})$. The resulting set of similarity vectors $\mathbf{s}_{\text{sim}}$ is clustered by GMM into two probabilistic sets:
@@ 2.3 Multi-Granularity Router
Existing methods rely on single predefined granularity for memory retrieval, limiting their ability to adaptively prioritize fine- or coarse-grained information based on query [19, 49, 42]. To address this, we propose an entropy-based router that adaptively selects the optimal granularity for each query.
[…]
$H^{g}=-\sum_{i=1}^{n}p^{g}_{i}(\boldsymbol{s}^{g})\log p^{g}_{i}(\boldsymbol{s}^{g}),\ \text{where}\ p^{g}_{i}(\boldsymbol{s}^{g})=\frac{\exp(\text{sim}(q,M_{i}^{g})/\lambda)}{\sum_{j=1}^{n}\exp(\text{sim}(q,M_{j}^{g})/\lambda)},\forall i\in\{1,\dots,n\}.$ (3)
where $H^{g}$ quantifies the uncertainty of matching $q$ to memories at granularity $g$. $\lambda$ is a hyperparameter to control the degree of entropy, and analyzed in Appendix D.
@@
where $G$ denotes the total number of granularities. This formulation ensures that granularities with lower entropy (indicating higher certainty) are assigned greater weights. Consequently, memories are reweighted to emphasize those associated with the query’s most confident granularities, eliminating the need for manual intervention.
@@ 2.4 Memory Retrieval and Filter
After constructing the multi-granularity memory associations and determining granularity weights, we retrieve relevant memories for a query $q$ by leveraging the graph-structured memory $\{M_{i},A_{i}\}\in\mathcal{M}\times\mathcal{A}$. To fully utilize inter-memory relationships, we employ the Personalized PageRank (PPR) algorithm [10] for context-aware ranking. The initial relevance score for each memory node $M_{i}$ is computed as a weighted aggregation across all granularities, using the router-assigned weights $w^{g}$ from Equation 4:
@@
Using the initial scores $\text{score}_{i}$ as personalized starting probabilities, we select the top $\alpha$ nodes as seed nodes and run the PPR algorithm on the memory association graph. This propagates relevance signals through the graph structure, emphasizing memories that are both directly relevant to the query and densely connected to other high-value memories. The final PPR scores rank memories by their global importance within the graph context. After convergence, we select the top-$K$ nodes as candidate contexts by their final PPR scores. The impact of $K$ is empirically analyzed in § 3.4.
LLM-Based Redundancy Filtering. To minimize noise and eliminate redundancy in the retrieved multi-granularity memories, we employ an LLM-based filtering mechanism. This mechanism processes the top-$K$ memories alongside the query $q$, using a carefully designed prompt (see Appendix H) to identify and discard irrelevant or repetitive content. As a result, the final context provided to the response generator is refined to focus exclusively on the most critical information relevant to $q$, ensuring a more concise and personalized response. We present case study on multi-granularity information in Appendix I.2 and case study on filtering in Appendix I.3.
@@ 3.1 Experimental Settings
Dataset and Metrics. Experiments are conducted on four comprehensive long-term memory datasets: LoCoMo-10 [30], Long-MT-Bench+ [33], LongMemEval-s [50], and LongMemEval-m [50], all focused on evaluating the capabilities of LLM agents in long-term conversations. Since our tasks require no training, the whole QA pairs in datasets are used as test set. Detailed dataset statistics are provided in Appendix A. To comprehensively evaluate model performance, we employ multiple metrics: F1 scores (as used by Maharana et al. [30]), BLEU, BERTScore, and ROUGE scores (following Pan et al. [33]). Additionally, we introduce GPT4o-as-Judge (GPT4o-J) [58], an evaluation setting where GPT4o assesses the alignment of a model’s response with the reference answer. The evaluation prompts are provided in Appendix H.
@@
Baselines. We compare MemGAS against various methods. (1) Full History: which utilizes all the latest conversation records, incorporating up to 128k tokens of context. Two strong retrieval models: (2) MPNet [43] and (3) Contriever [13]. Three memory-based conversational models: (4) LLM-Rsum [49], which uses LLMs to recursively summarize and update memory for contextually relevant responses; (5) MPC [19], which leverages pre-trained LLMs to create high-quality conversational agents; (6) SeCom [33], which segments memory into coherent topics and applies compression-based denoising to boost retrieval accuracy; Two structured RAG models: (7) HippoRAG 2 [8], which integrates knowledge graphs for efficient retrieval, and (8) RAPOTR [42], which enhances retrieval via recursive summary and hierarchical clustering into a tree structure. We also provide a comparison of different retrievers in Appendix C and error analysis in Appendix G.
@@
Implementation Details. For our experiments, we employ ‘gpt-4o-mini-2024-07-18’ as backbone for all tasks, including multi-granularity information generation and question-answer generation. To ensure fair comparisons, all baselines are implemented using consistent generation prompts. The temperature of LLMs is set to 0 to facilitate reproducibility. Additionally, a uniform top-3 document retrieval setting is applied across all models in the main results. We use Contriever [13] as the encoding model to generate embeddings for memory texts. The prompt templates are available in Appendix H. Detailed cost and efficiency are analyzed in Appendix F.

### v2

# From Single to Multi-Granularity: Toward Long-Term Memory Association and Selection of Conversational Agents

@@ 1 Introduction
Large Language Models (LLMs) have showcased remarkable conversational abilities, enabling them to serve as personalized assistants (Li et al., 2025a; Wang et al., 2024; Li et al., 2024b) and handle a wide range of applications, such as customer service (Pandya & Holia, 2023) and software development (Qian et al., 2024). However, as user-agent interactions increase, the volume of conversational history grows significantly. Despite their impressive conversational abilities, LLMs struggle with maintaining long-term conversational memory due to their limited context length (Liu et al., 2023b), which makes it challenging to retain a comprehensive record of user interactions and preferences over time. This limitation undermines their ability to generate coherent and personalized responses (Wu et al., 2025a; Pan et al., 2025; Xu et al., 2025; Li et al., 2025b; Chhikara et al., 2025). The development of retrieval-based external (non-parametric) memory systems (Packer et al., 2023; Park et al., 2023) has emerged as a promising solution. By storing interaction histories and retrieving relevant information when needed, memory systems enable LLMs to recall user-specific details from past dialogues and deliver more tailored responses.
@@
Recent studies on external memory systems of LLM agents primarily depend on the Retrieval-Augmented Generation (RAG) pipeline and have explored diverse aspects of memory segmentation and construction for effective retrieval (Wu et al., 2025b; Zhang et al., 2024a; Chen et al., 2025a; Chhikara et al., 2025; Zhong et al., 2024). For memory segmentation, existing approaches mainly adopt single-granularity to segment conversations. They utilize session-level chunks as retrieval units (Lu et al., 2023; Liu et al., 2023a; Team, 2023), while others employ finer-grained turn-level segmentation to capture details (Wu et al., 2025a; Yuan et al., 2023). Recent advances introduce topic-aware segmentation techniques (Pan et al., 2025; Tan et al., 2025) that group dialogue based on semantic coherence, enhancing topic-consistent retrieval. Additionally, several works (Wang et al., 2025; Team, 2023; Zhong et al., 2024; Sarthi et al., 2024) generate memory summary, condensing key information into compact representations to improve retrieval efficiency. Regarding memory construction, researchers have explored structured organization paradigms to enhance long-term knowledge retention (Li et al., 2024c; Edge et al., 2024; Guo et al., 2024). Methods like RAPTOR (Sarthi et al., 2024) and MemTree (Rezazadeh et al., 2024) employ hierarchical tree structures to encode multi-scale relations between memory units. Inspired by cognitive mechanisms (Teyler & DiScenna, 1986), HippoRAG (Gutiérrez et al., ; Gutiérrez et al., 2025) implement graph-based memory architectures that simulate neural memory consolidation processes, establishing relations between entities.
@@
Figure 1: An example of the multi-session query. The agent needs to synthesize information from both Conversation 1 and Conversation 2 to answer the question. The red underlines highlight information directly relevant to the answer, while the bold text emphasizes the same details between conversations. By generating multi-granularities (e.g., summary and keyword), the agent can better establish connections between conversations, enhancing full memory retrieval. Additionally, chart (a) illustrates the performance across different granularities, where ‘Suited Gran’ means the result using best-suited granularity for each query.
@@
However, despite their progress, current approaches exhibit two critical limits: i) Insufficient Multi-Granularity Memory Connection. While existing methods endeavor to organize memories into topological structures (e.g., knowledge graphs (Gutiérrez et al., 2025; Chhikara et al., 2025) or trees (Sarthi et al., 2024; Rezazadeh et al., 2024)), they predominantly concentrate on singular granularity levels—either entities or session summaries. This single-scale paradigm fails to model cross-granular interactions between memory units, resulting in retrieving only partial useful information. As demonstrated in Figure 1, answering multi-session queries requires establishing semantic links across granularities (e.g., connecting Conversation 1 and Conversation 2 through shared keyword/summary). Failure to establish links results in partial retrieval (e.g., retrieving only Conversation 1), which leads to incorrect answers. ii) Lack of Adaptive Multi-Granular Memory Selection. Current methods mainly rely on fixed granularity strategies (e.g., session/turn segmentation or LLM-generated summaries (Wang et al., 2025; Zhong et al., 2024; Wu et al., 2025a)), which often lead to incomplete context recall or noise due to improper granularity. Although topic-aware segmentation (Pan et al., 2025; Tan et al., 2025) enhances intra-chunk coherence, they lack adaptive mechanisms to select granularity for each query. Our empirical analysis in Figure 1(a) reveals that adaptively choosing best-suited granularity per query (e.g., balancing noise reduction in summaries/keywords with information retention in raw sessions) yields substantial performance gains. This highlights the necessity of granularity selection to resolve the inherent noise-information trade-off.
@@ 2 Methodology
This section defines the task and data format, constructs a dynamical memory association framework, details an entropy-based router for better-suited granularity, and outlines strategies for retrieving and filtering high-quality contextual information for response generation.
@@ 2.2 Multi-Granularity Association Construction
Existing methods mainly encode memories into vector libraries (e.g., session-level chunks) and directly retrieve information via similarity search (Izacard et al., 2021; Lee et al., 2023; Lu et al., 2023). However, such approaches overlook deeper associations between memories. To address this, we propose an associative memory construction process that captures multi-granular relationships.
@@
Dynamical Memory Association. When a new memory $\mathcal{M}_{\text{new}}$ is added, the current memory bank $\mathcal{M}_{\text{cur}}$ is updated as $\mathcal{M}_{\text{cur}}\leftarrow\mathcal{M}_{\text{cur}}\cup\mathcal{M}_{\text{new}}$. To establish associations between $\mathcal{M}_{\text{new}}$ and historical memories, a Gaussian Mixture Model (GMM)-based clustering strategy (Rasmussen, 1999) is used. All memories are encoded into dense vectors $e(\mathcal{M}_{\text{cur}})$ and $e(\mathcal{M}_{\text{new}})$ by Contriever (Izacard et al., 2021), and pairwise similarity scores are computed between $e(\mathcal{M}_{\text{new}})$ and each $e(\mathcal{M}_{j})\in e(\mathcal{M}_{\text{cur}})$. The resulting set of similarity vectors $\mathbf{s}_{\text{sim}}$ is clustered by GMM into two probabilistic sets:
@@ 2.3 Multi-Granularity Router
Existing methods rely on single predefined granularity for memory retrieval, limiting their ability to adaptively prioritize fine- or coarse-grained information based on query (Lee et al., 2023; Wang et al., 2025; Sarthi et al., 2024). To address this, we propose an entropy-based router that adaptively selects the better-suited granularity for each query.
Entropy-Driven Granularity Selection. For a query $q$, we first compute its similarity to all memory chunks across each granularity level $g\in\{\text{session, turn, summary, keyword}\}$. Let $\bm{s}^{g}=[\text{sim}(q,M_{1}^{g}),\dots,\text{sim}(q,M_{n}^{g})]$ denote the similarity scores between $q$ and n memories chunks at granularity $g$. We normalize $\bm{s}^{g}$ into a probability distribution $p^{g}(\bm{s}^{g})$ and calculate its Shannon entropy:
$H^{g}=-\sum_{i=1}^{n}p^{g}_{i}(\bm{s}^{g})\log p^{g}_{i}(\bm{s}^{g}),\ \text{where}\ p^{g}_{i}(\bm{s}^{g})=\frac{\exp(\text{sim}(q,M_{i}^{g})/\lambda)}{\sum_{j=1}^{n}\exp(\text{sim}(q,M_{j}^{g})/\lambda)},\forall i\in\{1,\dots,n\}.$ (3)
where $H^{g}$ quantifies the uncertainty of matching $q$ to memories at granularity $g$. $\lambda$ is a hyperparameter to control the degree of entropy, and analyzed in Appendix F.
@@
where $G$ denotes the total number of granularities. This formulation ensures that granularities with lower entropy (indicating higher certainty) are assigned greater weights. Consequently, memories are reweighted to emphasize those associated with the query’s most confident granularities, eliminating the need for manual intervention. We present a theoretical analysis of granularity association and routers in Appendix H.
@@ 2.4 Memory Retrieval and Filter
After constructing the multi-granularity memory associations and determining granularity weights, we retrieve relevant memories for a query $q$ by leveraging the graph-structured memory $\{M_{i},A_{i}\}\in\mathcal{M}\times\mathcal{A}$. To fully utilize inter-memory relationships, we employ the Personalized PageRank (PPR) algorithm (Haveliwala, 2002) for context-aware ranking. The initial relevance score for each memory node $M_{i}$ is computed as a weighted aggregation across all granularities, using the router-assigned weights $w^{g}$ from Equation 4:
@@
Using the initial scores $\text{score}_{i}$ as personalized starting probabilities, we select the top $\alpha$ nodes as seed nodes and run the PPR algorithm on the memory association graph. This propagates relevance signals through the graph structure, emphasizing memories that are both directly relevant to the query and densely connected to other high-value memories. The final PPR scores rank memories by their global importance within the graph context. After convergence, we select the top-$K$ nodes as candidate contexts by their final PPR scores. The impact of $K$ is empirically analyzed in Section § 3.4.
LLM-Based Redundancy Filtering. To minimize noise and eliminate redundancy in the retrieved multi-granularity memories, we employ an LLM-based filtering mechanism. This mechanism processes the top-$K$ memories alongside the query $q$, using a carefully designed prompt (see Appendix I) to identify and discard irrelevant or repetitive content. As a result, the final context provided to the response generator is refined to focus exclusively on the most critical information relevant to $q$, ensuring a more concise and personalized response. We present case study on multi-granularity information in Appendix J.2 and case study on filtering in Appendix J.3.
@@ 3.1 Experimental Settings
Dataset and Metrics. Experiments are conducted on four comprehensive long-term memory datasets: LoCoMo (Maharana et al., 2024), Long-MT-Bench+ (Pan et al., 2025), LongMemEval-s (Wu et al., 2025a), and LongMemEval-m (Wu et al., 2025a), all focused on evaluating the capabilities of LLM agents in long-term conversations. Since our task is training-free, the whole QA pairs in datasets are used for evaluation. Detailed dataset statistics are provided in Appendix A. To comprehensively evaluate model performance, we employ multiple metrics: F1 scores (as used by Maharana et al. (2024)), BLEU (with 4-gram by default), BERTScore, and ROUGE scores (following Pan et al. (2025)). Additionally, we introduce GPT4o-as-Judge (GPT4o-J) (Zheng et al., 2023), an evaluation setting where GPT4o assesses the alignment of a model’s response with the reference answer. The evaluation prompts are provided in Appendix I.
@@
Baselines. We compare MemGAS against various methods. (1) Full History: which utilizes all the latest conversation records, incorporating up to 128k tokens of context. Two strong retrieval models: (2) MPNet (Song et al., 2020) and (3) Contriever (Izacard et al., 2021). Four memory-based conversational models: (4) RecurSum (Wang et al., 2025), which uses LLMs to recursively summarize and update memory for contextually relevant responses; (5) MPC (Lee et al., 2023), which composes LLM with prompting and external memory; (6) A-Mem (Xu et al., 2025), which organizes memories through generating note and links; (7) SeCom (Pan et al., 2025), which segments memory into coherent topics and applies compression-based denoising to boost retrieval; Two structured RAG models: (8) HippoRAG 2 (Gutiérrez et al., ), which integrates knowledge graphs for efficient retrieval, and (9) RAPOTR (Sarthi et al., 2024), which enhances retrieval via recursive summary and hierarchical clustering into a tree structure. More details can be found in Appendix B.
Table 1: QA Performance. Contriever is used as the retrieval backbone (except for Full History and MPNet), with GPT4o-mini as the generator. Bolded values indicate the best results, while underlined values represent the second-best results. All metrics are the larger, the better. ‘4o-J’ denotes GPT4o-as-Judge, ‘B-4’ stands BLEU4, ‘R-n’ denotes Rouge-n and ‘BS’ stands BERTScore. The unit of latency is seconds.
Model 4o-J F1 B-4 R-1 R-2 R-L BS Avg. Tokens Avg. Latency
LongMemEval-s
Full History 50.60 11.48 1.40 12.10 5.47 10.85 83.07 103,137 9.39
MPNet (Song et al., 2020) 53.20 13.96 2.21 14.49 6.78 12.93 83.72 8,173 1.82
Contriever (Izacard et al., 2021) 55.40 13.78 2.21 14.46 6.93 12.89 83.70 8,286 1.85
MPC (Lee et al., 2023) 53.80 13.60 1.74 14.27 6.49 12.95 83.49 8,457 2.66
RecurSum (Wang et al., 2025) 35.40 12.29 2.09 13.01 5.55 11.52 83.60 8,853 1.94
SeCom (Pan et al., 2025) 56.00 12.95 2.25 13.80 6.09 11.93 83.51 2,741 1.67
HippoRAG 2(Gutiérrez et al., 2025) 57.60 14.73 2.15 15.30 7.36 13.83 83.86 8,530 4.51
RAPTOR (Sarthi et al., 2024) 32.20 12.08 1.90 12.73 5.82 11.25 83.50 6,254 2.25
A-Mem (Xu et al., 2025) 55.60 13.73 2.11 14.82 6.81 12.98 83.88 9,018 2.59
MemGAS (Ours) 60.20 20.38 4.22 21.05 10.47 19.47 85.21 8,829 2.55
LongMemEval-m
Full History 12.20 5.70 0.78 6.27 2.08 5.28 81.62 128,000 12.88
MPNet (Song et al., 2020) 37.80 10.76 1.70 11.46 4.76 10.03 83.09 8,352 1.83
Contriever (Izacard et al., 2021) 42.80 11.88 1.66 12.56 5.63 11.02 83.28 8,467 1.92
MPC (Lee et al., 2023) 37.80 11.28 1.37 11.93 5.12 10.57 82.98 8,428 2.75
RecurSum (Wang et al., 2025) 23.80 10.04 1.70 10.89 4.26 9.21 83.12 8,927 1.99
SeCom (Pan et al., 2025) 42.80 11.33 1.79 12.03 5.07 10.49 83.36 2,821 1.63
MemGAS (Ours) 45.40 16.85 3.39 17.60 8.25 16.14 84.69 8,852 2.45
LoCoMo
Full History 33.43 12.23 1.84 12.70 5.66 11.73 84.07 20,078 4.92
MPNet (Song et al., 2020) 38.07 14.44 2.35 14.90 6.83 13.90 84.42 2,472 1.29
Contriever (Izacard et al., 2021) 40.33 15.66 2.67 16.01 7.68 15.00 84.65 2,348 1.24
MPC (Lee et al., 2023) 40.38 14.81 1.99 15.10 6.83 14.13 84.43 2,683 1.95
RecurSum (Wang et al., 2025) 22.56 9.14 0.99 9.82 3.38 8.98 83.45 3,074 1.58
SeCom (Pan et al., 2025) 44.21 13.79 2.30 14.28 6.17 13.30 84.04 1,021 1.02
HippoRAG 2(Gutiérrez et al., 2025) 45.62 16.66 2.91 17.01 8.27 15.93 84.88 2,991 3.56
RAPTOR (Sarthi et al., 2024) 31.72 14.55 2.88 15.09 7.49 14.18 84.48 1,931 1.73
A-Mem (Xu et al., 2025) 40.81 14.72 2.83 16.22 7.71 14.89 84.72 3,042 1.98
MemGAS (Ours) 41.07 17.66 3.61 18.00 8.93 16.99 85.13 2,825 1.88
LongMTBench+
Full History 67.44 36.07 11.32 37.90 20.51 29.25 87.81 19,194 4.72
MPNet (Song et al., 2020) 63.89 36.09 11.26 38.39 20.58 28.86 87.84 12,143 3.21
Contriever (Izacard et al., 2021) 63.54 36.30 11.59 38.17 21.65 29.67 87.90 11,941 3.27
MPC (Lee et al., 2023) 61.81 31.52 7.97 33.56 17.27 25.20 86.51 12,289 3.98
RecurSum (Wang et al., 2025) 24.65 26.58 6.91 29.23 11.93 20.90 86.11 13,527 3.41
SeCom (Pan et al., 2025) 64.58 36.68 12.01 38.81 21.44 29.65 87.88 4,714 2.68
HippoRAG 2(Gutiérrez et al., 2025) 63.54 35.64 11.05 37.61 20.37 28.76 87.70 13,583 6.14
RAPTOR (Sarthi et al., 2024) 59.72 37.69 13.47 40.08 21.68 30.88 88.38 10,631 3.47
A-Mem (Xu et al., 2025) 65.73 36.82 11.36 38.92 20.88 29.14 87.92 13,735 3.95
MemGAS (Ours) 69.44 41.49 15.62 43.69 24.45 34.66 88.96 12,873 3.85
@@
Implementation Details. We use ‘gpt-4o-mini-2024-07-18’ as the backbone for all tasks, including multi-granularity information generation and QA. To ensure fairness, all baselines share consistent generation prompts. The temperature of LLMs is set to 0 for reproducibility, and all models operate in a zero-shot setting with prompt templates detailed in Appendix I. A consistent top-3 session retrieval setting is applied across all models for fair comparison, with top-3 segments used for SeCom and sessions/summaries for RAPTOR. Contriever is used as the encoding model to generate embeddings for memory texts. LongMTBench+ is excluded due to the lack of a ground-truth of retrieval. RAPTOR and A-Mem retrieval cannot be evaluated as their retrieved information cannot be judged. Results for RAPTOR, A-Mem, and HippoRAG on LongMemEval-m are unavailable due to high runtime. We compared various methods across different retrievers, generators, and query types, as detailed in Appendix E.1 E.2 and E.3. Additionally, we provide a hyperparameter analysis in Appendix F and an error analysis in Appendix G. The additional cost and efficiency of the methods are discussed in Appendix D.